Please do by hand. Thanks in advance.

2. Let X and Y be two random variables. If Var(X) = 4, Var(Y) = 16, and Cov(X,Y) = 2, then what is Var(3Y - 2x)?
Answer #1

By the formula:

Var(aX - bY) = a2 Var(X) + b2 Var(Y) - 2*a*b*Cov(X, Y)

Hence,

Var(3Y - 2X) = 9*Var(Y) + 4*Var(X) - 2*3*2*Cov(X, Y) = 9*16 + 4*4 - 2*3*2*2 = 136